# LADICA: A Large Shared Display Interface for Generative AI Cognitive Assistance in Co-Located Team Collaboration

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Interactive systems and tools
@@ 1. INTRODUCTION
In this paper, we aim to design and develop an intelligent large shared display interface that incorporates LLM assistance into human-led co-located group collaboration. The primary goal is to provide dynamic cognitive support that enhances and supports beneficial practices in human-human interactions. In a formative study aimed at understanding how teams use whiteboards during in-person collaboration, we found that coworkers frequently rely on a shared whiteboard for brainstorming, grouping analysis and grounding their verbal discussions. This study pinpointed several key challenges participants face in such collaborations, as outlined Section 3.2. Through conducting scenario-based design workshops, we gathered insights into users’ concerns about the AI dominance and the potential over-reliance on AI in group collaborations. From these insights, we identified four specific design goals: (1) facilitate the cognitive process of task breakdown and diverse ideation; (2) enhance collaboration by fostering mutual awareness; (3) offer support for comprehensive analysis from various perspectives and dimensions; (4) encourage synchronization among team members of discussion entities and workspace activities.
@@
  • The development of LADICA, a large shared display interface that utilizes LLMs to provide multifaceted cognitive assistance using a three-layer approach in co-located group collaboration.
@@ 2.2. Team Cognition
Team cognition is central to team effectiveness. Warner et al. (Warner et al., 2005) proposed a cognitive model of the team cognition consisting processs at two levels: Meta-cognition and Macro-cognition.
@@ 3.1.1. Participants
We recruited 8 participants (5 males, 3 females) with experience in in-person team collaboration from a university’s mailing list. Each participant was compensated with a 15 USD gift card. Our participants consisted of four undergraduate and four graduate students, with majors in computer science, business, applied math, and social science. They had participated in various forms of in-person teamwork, including planning club events, discussing course projects, organizing trips, and planning software development. They all had previously used a shared physical whiteboard for co-located teamwork and had experience with digital tools such as Excel, Notion, Miro, and Microsoft OneNote for group work. Seven of them had experience in using LLMs such as ChatGPT and Gemini.
@@ 3.1.2. Procedure
Our study involved two in-person sessions, each with four participants, lasting one hour. These sessions were divided into a 30-minute focus group and a 30-minute scenario-based design workshop [^2] The study protocol has been reviewed and approved by the IRB at our institution.. In focus groups, participants discussed their past experiences, challenges, and unmet needs when working with a co-located team. Additionally, we also asked them to discuss their concerns about AI intervention during team collaboration.
Following the focus group, we presented a scenario of using an AI-enhanced whiteboard to discuss and plan a spring break trip with friends. Participants were invited to illustrate their ideal AI features on paper storyboards or write down their ideas on sticky notes. The study sessions were recorded, transcribed, and analyzed using the reflexive thematic analysis method (Braun and Clarke, 2012).
@@ 3.2.1. Whiteboard is used for supporting ideation, analysis and conversation
In our formative study, we found that in-person teams primarily used a shared whiteboard for three key purposes: visually expressing individual ideas, collectively understanding and analyzing tasks, and providing a visual anchor for discussions. Visualization helps align members’ mental models, aiding in avoiding knowledge gaps in verbal discussions” (P3) andreducing the cognitive load of remembering shared information” (P6). Additionally, as teams often focus on ideation and planning stages in meetings (P3, P4, P5, P7, P8), the whiteboard supports both bottom-up and top-down approaches to developing goals. For example, some teams use it for unstructured ideation to organize diverse ideas meaningfully” (P7), while others prefer a structured approach tocreate a top-down plan and then delegate tasks” (P4). Whiteboards also help teams base discussions on previously visualized information, allowing them to “refer back to clarify which ideas are being discussed and identify new ideas based on the existing ones” (P2). Based on these insights, we are developing a digital whiteboard system designed to enhance cognitive processes during group ideation, analysis, and conversation, and to support flexible collaboration styles.
@@
Consistent with prior research (Ez-Zaouia and Carrillo, 2023; Dijkstra et al., 2008; Drapeau, 2014), our participants highlighted cognitive and social-motivational challenges in in-person collaboration. Cognitively, individuals often struggle with divergent and critical thinking, finding it difficult to quickly generate a wide range of ideas or comment on others’ ideas” (P6) and toarticulate thoughts relevantly and clearly” (P4). Socially, the presence of team members can create pressures that reduce effectiveness, such as conformity pressure leading to narrow idea selection (P3, P5) and fear of judgment inhibiting idea sharing (P1, P3, P7). Additionally, teams may find it challenging to understand task structure and prioritize topics and subtasks during brainstorming (P3, P8). To address these issues, participants suggested integrating AI features to inspire and refine ideas, facilitate expression, and clarify discussion points and task structures, aiming to enhance both the creative and organizational facets of team collaboration.
@@ 3.2.3. Maintaining mutual awareness during parallel team work is crucial yet difficult
Mutual awareness is crucial for managing interdependent tasks and overall progress in teams (Liu et al., 2016; Tang et al., 2006). Our study identified challenges in maintaining this awareness during parallel tasks within a shared workspace. For instance, P5 noted, when working on different parts of a task on a whiteboard, it’s challenging to remain mindful of how my work relates to and impacts others and vice versa. This awareness is crucial for us to coordinate interdependent and parallel tasks.” P7 added,we tend to miss collaboration opportunities when people focus on different parts without realizing the connection among each other’s ongoing work.” To mitigate these issues, participants recommended integrating AI to highlight connections between individual contributions, enhancing understanding of how ideas interrelate and revealing opportunities for dialogue.
@@ 3.2.4. The need for varied analytical perspectives in organizing and comparing information
Participants noted that whiteboards are crucial for organizing group ideas into themes, fostering a bottom-up understanding, stimulating informal discussions, and identifying goal-related gaps. They also expressed a need for diverse analytical perspectives to organize information and track updates, essential for revealing common themes and patterns in their proposals (P6). However, the current digital whiteboard tools’ limited support for flexible information grouping and tracking perspective evolution restricts experimenting with various organizational methods and gaining deeper insights. Some participants (P3, P5) also desired the ability to save and reload different workspace snapshots, facilitating experimentation with various groupings and explorations. In response, during design workshops, participants envisioned scenarios where AI could suggest different categorization strategies for team-generated ideas and automatically group them, enhancing the exploration of various analytical perspectives.
@@ 3.2.5. Challenges in Synchronizing and Anchoring Discussion on Shared Content
In co-located group collaboration, discussions are typically informal, active, and fluid, which enhances team cohesion and engagement. However, participants noted challenges in maintaining synchronization between the discussion and the content on the whiteboard (P2, P5, P8). For example, P3 stated, ”It’s easy for interesting points to be lost when everyone is engaged in the conversation and no one is documenting them on the whiteboard.” Moreover, when the whiteboard is overloaded with information, it becomes difficult for teams to quickly locate and concentrate on details relevant to the ongoing conversation (P3, P4, P7), which is crucial for anchoring discussions and refreshing shared understanding (P3, P4). To address these issues, participants recommended developing features that could automatically extract key points from discussions and emphasize elements on the whiteboard pertinent to the current conversation.
@@ 3.2.6. Concerns regarding AI involvement in group collaboration
While participants appreciated the AI features designed to enhance team collaboration, they also voiced several concerns about AI involvement. There was worry that AI might dominate the natural dynamics of human collaboration, with fears that AI-driven interactions could overshadow human-to-human interactions (P1, P3, P4, P6, P7). As P5 mentioned, ”I have little worry about AI taking over control of discussions, though I feel like reasonable humans still know what they want from the discussion.” Additionally, there were concerns that relying on AI for idea generation and planning could reduce critical thinking and creativity among team members (P2, P4, P5, P8). Privacy issues were also raised, with a preference expressed for AI assistance to be optional and only activated upon explicit request (P1, P4, P7).
@@ 3.3. Design Goals
  • DG1: Enhancing dynamic ideation and knowledge sharing within groups.
  • DG2: Fostering collaborative cues without overriding human interaction.
  • DG3: Enabling diverse and comprehensive perspectives in analyzing shared content.
  • DG4: Streamlining discussion and workspace synchronization.
@@ 4.1. The Three-Layer Structure for Cognitive Assistance
The first layer, called the idea repository, focuses on the initial stage of group ideation, where team members contribute their individual ideas, knowledge, and opinions regarding the team’s goals. This layer supports the team by organizing these individual contributions in a way that everyone can see and build upon, encouraging the expansion of ideas and a deeper understanding of the task at hand. Additionally, it helps the team decompose the overall goal into smaller, manageable tasks, assisting the team’s meta-cognition of understanding and monitoring the group task structure.
The second layer, the affinity lens, aims to facilitate the organization of shared ideas. This layer focuses on assisting team’s cognitive process of comparing and grouping ideas from various perspectives to identify structured themes within their ideas. At the meta-cognitive level, this layer offers hierarchical affinity analysis, providing prompts that encourage hierarchical thinking. At the macro-cognitive level, this layer assists in the exploratory grouping process by offering hints for different analytical perspectives, automating multi-faceted affinity grouping, and allowing for the rollback to previously saved snapshots.
The final layer, known as the discussion reference, is designed to foster group discussions by highlighting the connections between different ideas. To create discussion opportunity among teams, it displays hints about the relationships among team ideas to support the creation and maintenance of mutual awareness at the meta-cognitive level. For macro-cognitive assistance in group discussions, LADICA facilitates the connection between past and ongoing discussions by enabling the retrieval of relevant ideas on-screen during live discussions. Furthermore, to promote collective thinking, it provides group-based discussion hints as teams focus on particular thematic groups.
@@ 4.3.1. Idea repository layer
Figure 6. The interaction flow of discussion-based key information extraction and retrieval. Users can start recording the ongoing discussion by pressing “Start Recording” under the “Speech” menu. The live transcription is displayed simultaneously. By pressing “Get relevant ideas,” LADICA will identify existing ideas related to the discussion. By pressing “Extract key information,” LADICA will extract key information from the discussion and also highlight related existing idea notes.
@@
During the ideation stage, co-located teams typically share their ideas verbally, which can add extra cognitive load when they need to remember and transfer the key information from their discussion onto an external shared workspace (DG4). To alleviate this inconvenience, LADICA offers a discussion-based key information extraction feature that allows teams to focus on discussing ideas while automatically highlighting key information once the discussion ends. As shown in Figure 6, teams can press the “Speech” button on the workspace menu and select “Start recording” to record their discussion. LADICA will display the latest transcription of their conversation as it progresses. When the team concludes a section of their discussion, they can press “Extract key information” to identify the key points from their dialogue. If a key point relates to existing idea notes on screen, LADICA will also show the related note in each information card. Teams can then press on a key information card to create an idea note containing the same content.
@@ 4.3.3. Discussion reference layer
Establishing and preserving mutual awareness is essential for team coordination and recognizing opportunities for collaboration. To support this meta-cognitive process, LADICA is able to provide relation hints among ideas and update them periodically throughout the collaboration (DG2). As shown in Figure 9, users can press “Action” then “Toggle relation hints” in the workspace menu to activate the relation hint generation. In addition, users can toggle the “cross-user only” to only generate ideas created by different team members. To reflect the relations among the latest ideas, LADICA will update the relation hints every 10 seconds when the function is activated. Note that this feature is designed to indicate relationships to catalyze initial discussion rather than to suggest specific discussion content. To achieve this, we instruct the LLM to generate relation types based on the ConceptNet relation taxonomy (Speer et al., 2017), which encompasses a wide range of types and maintains a level of granularity that provides teams with ample room to explore the relationships.
@@
Our findings indicate that team members often struggle to keep the whiteboard content aligned with the dynamic flow of their discussions. To address this challenge, LADICA is designed to intelligently retrieve relevant existing idea nodes based on the content of the conversation (DG4). Users can activate this feature by pressing the “Start Recording” button at the beginning of their collaborative session (Figure 6). This prompts LADICA to start recording while simultaneously analyzing the ongoing discussion to identify mentions or references related to existing ideas. Upon ending the session with the ’Stop Recording’ button, LADICA processes the recorded conversation. The system then retrieves the existing ideas that have relevance score with its corresponding dialogue reference in the discussion over a predefined threshold (see Section 4.4.5). For each relevant ideas, LADICA also shows the corresponding dialogue reference.
@@
A crucial opportunity for generating new insights is to reflect on team’s ideas within group, which sparks diverse perspective in discussion. To facilitate this process, LADICA provides group-based discussion hint (Figure 10) that is able to generate hints of discussion topics about the theme of the group and existing ideas within the group. Teams can access this feature by pressing [Uncaptioned image] in the group sidebar (Figure 1, D). Users can provide further instructions for hint generation. It is important to note that the hints are not intended to generate ideas directly; instead, they are designed to inspire potential new directions for team thinking and discussion (DG2).

### v2

@@ Abstract.
journalyear: 2025copyright: ccconference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3713289isbn: 979-8-4007-1394-1/25/04ccs: Human-centered computing Interactive systems and tools
@@ 1. INTRODUCTION
In this paper, we aim to design and develop an intelligent large shared display interface that incorporates LLM assistance into human-led co-located group collaboration for brainstorming and sensemaking activities. The primary goal is to provide dynamic cognitive support that enhances and supports beneficial practices in human-human interactions. In a formative study aimed at understanding how teams use whiteboards during in-person collaboration, we found that coworkers frequently rely on a shared whiteboard for brainstorming, grouping analysis and grounding their verbal discussions. This study pinpointed several key challenges participants face in such collaborations, as outlined Section 3.2. Through conducting scenario-based design workshops, we gathered insights into users’ concerns about the AI dominance and the potential over-reliance on AI in group collaborations. From these insights, we identified four specific design goals: (1) facilitate the cognitive process of task breakdown and diverse ideation; (2) enhance collaboration by fostering mutual awareness; (3) offer support for comprehensive analysis from various perspectives and dimensions; (4) encourage synchronization among team members of discussion entities and workspace activities.
@@
  • The development of LADICA, a large shared display interface that utilizes LLMs to provide multifaceted cognitive assistance using a three-layer approach in co-located group collaboration for brainstorming and sensemaking activities.
@@ 2.2. Team Cognition
Team cognition is central to team effectiveness. Warner et al. (Warner et al., 2005) proposed a cognitive model of the team cognition consisting process at two levels: Meta-cognition and Macro-cognition.
@@ 3.1.1. Participants
We recruited eight participants (five males, three females) with experience in in-person team collaboration from a university’s mailing list. Each participant was compensated with a \$15 USD gift card. Our participants consisted of four undergraduate and four graduate students, with majors in computer science, business, applied math, and social science. They had participated in various forms of in-person teamwork, including planning club events, discussing course projects, organizing trips, and planning software development. Seven of them have used AI products (e.g. ChatGPT, Gemini) in the past. One participant had little to no knowledge of AI, four had a basic understanding of AI concepts, and three had experience developing AI systems or possessed a strong understanding of AI. Besides, they all had previously used a shared physical whiteboard for collaborative teamwork and had experience with digital tools such as Excel, Notion, Miro, and Microsoft OneNote for group work.
@@ 3.1.2. Procedure
Our study involved two in-person sessions, each with four participants, lasting one hour. Two authors attended each session to facilitate the study and take notes. These sessions were divided into a 30-minute focus group and a 30-minute scenario-based design workshop [^2] The study protocol has been reviewed and approved by the IRB at our institution.. In focus groups, participants discussed their past experiences, challenges, and unmet needs when working with a co-located team. Additionally, we also asked them to discuss their concerns about AI intervention during team collaboration.
Following the focus group, we presented a scenario of using an AI-enhanced whiteboard to discuss and plan a spring break trip with friends. Participants were invited to illustrate their ideal AI features on paper storyboards or write down their ideas on sticky notes. The study sessions were recorded using an external video camera, then transcribed and analyzed using the reflexive thematic analysis method (Braun and Clarke, 2012).
@@ 3.2.1. Whiteboard is used for supporting ideation, analysis and conversation
In our formative study, we found that in-person teams primarily used a shared whiteboard for three key purposes: visually expressing individual ideas, collectively understanding and analyzing tasks, and providing a visual anchor for discussions. Through the focus group discussion, participants have agreed that visualization helps align members’ mental models, aiding in “avoiding knowledge gaps in verbal discussions” (P3) and “reducing the cognitive load of remembering shared information” (P6). Additionally, as teams often focus on ideation and planning stages in meetings (P3, P4, P5, P7, P8), the whiteboard supports both bottom-up and top-down approaches to developing goals. For example, during design workshops, P7 brainstormed features of meaningfully organizing diverse ideas in open-ended ideation, while P4 sketched a more structured approach to create a top-down plan and then delegate tasks . Whiteboards also help teams base discussions on previously visualized information, allowing them to “refer back to clarify which ideas are being discussed and identify new ideas based on the existing ones” (P2). Based on these insights, we are developing a digital whiteboard system designed to enhance cognitive processes during group ideation, analysis, and conversation, and to support flexible collaboration styles.
@@
Consistent with prior research (Ez-Zaouia and Carrillo, 2023; Dijkstra et al., 2008; Drapeau, 2014), our participants highlighted cognitive and socially motivational challenges during in-person collaboration. The focus groups showed that individuals often struggle with divergent and critical thinking, finding it difficult to “quickly generate a wide range of ideas or comment on others’ ideas” (P6) and to “articulate thoughts relevantly and clearly” (P4). Socially, the presence of team members can create pressures that reduce effectiveness, such as conformity pressure leading to narrow idea selection (P3, P5) and fear of judgment inhibiting idea sharing (P1, P3, P7). Additionally, teams may find it challenging to understand task structure and prioritize topics and subtasks during brainstorming (P3, P8). To address these issues, participants suggested integrating AI features to inspire and refine ideas, facilitate expression, and clarify discussion points and task structures, aiming to enhance both the creative and organizational facets of team collaboration.
@@ 3.2.3. Maintaining mutual awareness during parallel team work is crucial yet difficult
Mutual awareness is crucial for managing interdependent tasks and overall progress in teams (Liu et al., 2016; Tang et al., 2006). The focus groups identified challenges in maintaining this awareness during parallel tasks within a shared workspace. For instance, P5 noted, “when working on different parts of a task on a whiteboard, it’s challenging to remain mindful of how my work relates to and impacts others and vice versa. This awareness is crucial for us to coordinate interdependent and parallel tasks.” P7 added,“we tend to miss collaboration opportunities when people focus on different parts without realizing the connection among each other’s ongoing work.” To mitigate these issues, participants recommended integrating AI to highlight connections between individual contributions, enhancing understanding of how ideas interrelate and revealing opportunities for dialogue.
@@ 3.2.4. The need for varied analytical perspectives in organizing and comparing information
Participants noted that whiteboards are crucial for organizing group ideas into themes, fostering a bottom-up understanding, stimulating informal discussions, and identifying goal-related gaps. They also expressed a need for diverse analytical perspectives to organize information and track updates, essential for revealing common themes and patterns in their proposals (P6). However, the current digital whiteboard tools’ limited support for flexible information grouping and tracking perspective evolution restricts experimenting with various organizational methods and gaining deeper insights. In the design workshops, some participants (P3, P5) sketched features for saving and reloading different workspace snapshots, facilitating experimentation with various groupings and explorations. In response, during design workshops, participants envisioned scenarios where AI could suggest different categorization strategies for team-generated ideas and automatically group them, enhancing the exploration of various analytical perspectives.
@@ 3.2.5. Challenges in Synchronizing and Anchoring Discussion on Shared Content
In co-located group collaboration, discussions are typically informal, active, and fluid, which enhances team cohesion and engagement. However, participants noted challenges in maintaining synchronization between the discussion and the content on the whiteboard (P2, P5, P8). For example, P3 stated, “It’s easy for interesting points to be lost when everyone is engaged in the conversation and no one is documenting them on the whiteboard.” Moreover, when the whiteboard is overloaded with information, it becomes difficult for teams to quickly locate and concentrate on details relevant to the ongoing conversation (P3, P4, P7), which is crucial for anchoring discussions and refreshing shared understanding (P3, P4). To address these issues, participants recommended developing features that could automatically extract key points from discussions and emphasize elements on the whiteboard pertinent to the current conversation.
@@ 3.2.6. Concerns regarding AI involvement in group collaboration
While participants appreciated the AI features designed to enhance team collaboration, they also voiced several concerns about AI involvement. There was worry that AI might dominate the natural dynamics of human collaboration, with fears that AI-driven interactions could overshadow human-to-human interactions (P1, P3, P4, P6, P7). As P5 mentioned, “I have little worry about AI taking over control of discussions, though I feel like reasonable humans still know what they want from the discussion.” Additionally, there were concerns that relying on AI for idea generation and planning could reduce critical thinking and creativity among team members (P2, P4, P5, P8). Privacy issues were also raised, with a preference expressed for AI assistance to be optional and only activated upon explicit request (P1, P4, P7).
@@ 3.3. Design Goals
  • DG1: Enhancing dynamic ideation and knowledge sharing within groups. Social and motivational barriers, such as fear of judgment and social loafing, can prevent team members from sharing ideas freely, affecting inclusivity and interaction quality. Cognitive theories and our formative study highlight the need for metacognitive support to help teams understand complex tasks. Team members also need help articulating ideas, refining them during discussions, and recognizing overlooked ideas. To address these challenges, our proposed system provides cognitive assistance to support ideation and promote knowledge sharing.
  • DG2: Streamlining bidirectional discussion and work space synchronization. Maintaining mutual awareness is crucial for identifying collaborative opportunities, especially when team members work on different tasks simultaneously. Our formative study shows that this awareness can be difficult to achieve due to cognitive overload. While participants are open to AI assistance, they worry that too much AI involvement could disrupt natural team dynamics and reduce direct interaction. Therefore, our system is designed to help teams see how their contributions are connected, without dominating the collaboration process. It avoids automatically completing tasks that require group discussion or overly controlling the discussion flow or sequence.
  • DG3: Enabling diverse and comprehensive perspectives in analyzing shared content. Our formative study showed that teams need to analyze and group ideas from different perspectives to uncover themes and connections, but this process can be challenging. Teams also need to track how these patterns evolve during collaboration and have the ability to save and reload workspace snapshots to experiment freely. To meet these needs, our system should support grouping analysis from multiple perspectives, helping teams in both the creation and discussion phases.
  • DG4: Fostering collaborative cues without overriding human interaction. Co-located group collaboration often involves a lot of informal discussion. Our formative study found that synchronizing these discussions with the shared workspace can be cognitively challenging. To address this, our system identifies key points from discussions and highlights relevant information already in the workspace. This helps bridge the gap between verbal exchanges and visual representations, ensuring that important discussion elements are effectively integrated.
@@ 4.1. The Three-Layer Structure for Cognitive Assistance
The first layer, called the idea repository, focuses on the initial stage of group ideation, where team members contribute their individual ideas, knowledge, and opinions regarding the team’s goals. To enhance dynamic ideation and knowledge sharing (DG1), this layer supports the team by organizing these individual contributions in a way that everyone can see and build upon, encouraging the expansion of ideas and a deeper understanding of the task at hand. Additionally, it helps the team decompose the overall goal into smaller, manageable tasks, assisting the team’s meta-cognition of understanding and monitoring the group task structure. Besides, this layer helps synchronize and offload verbal ideas to the shared workspace during the ideation phase (DG2).
The second layer, the affinity lens, aims to facilitate the organization of shared ideas. This layer focuses on assisting team’s cognitive process of comparing and grouping ideas from various perspectives to identify structured themes within their ideas (DG3). At the meta-cognitive level, this layer offers hierarchical affinity analysis, providing prompts that encourage hierarchical thinking. At the macro-cognitive level, this layer assists in the exploratory grouping process by offering hints for different analytical perspectives, automating multi-faceted affinity grouping, and allowing for the rollback to previously saved snapshots.
The final layer, known as the discussion reference, is designed to foster group discussions by highlighting the connections between different ideas. To create discussion opportunity among teams, it displays hints (DG4) about the relationships among team ideas to support the creation and maintenance of mutual awareness at the meta-cognitive level. For macro-cognitive assistance in group discussions, LADICA facilitates the connection between past and ongoing discussions by enabling the retrieval of relevant ideas on-screen during live discussions, synchronizing the conversation with the content in the shared workspace (DG2). Furthermore, to promote collective thinking, it provides group-based discussion hints as teams focus on particular thematic groups.
@@ 4.3.1. Idea repository layer
Figure 6. The interaction flow of discussion-based key information extraction and retrieval. Users can start recording the ongoing discussion by pressing “Start Recording” under the “Speech” menu. The live transcription is displayed simultaneously. By pressing “Get relevant ideas,” LADICA will identify existing ideas related to the discussion, where each card on the right showing an existing idea relevant to the ongoing discussion and corresponding transcription; By pressing “Extract key information”, LADICA will extract key information from the discussion and also highlight related existing idea notes, as shown in cards on the right in the lower figure.
@@
During the ideation stage, co-located teams typically share their ideas verbally, which can add extra cognitive load when they need to remember and transfer the key information from their discussion onto an external shared workspace (DG2). To alleviate this inconvenience, LADICA offers a discussion-based key information extraction feature that allows teams to focus on discussing ideas while automatically highlighting key information once the discussion ends. As shown in Figure 6, teams can press the “Speech” button on the workspace menu and select “Start recording” to record their discussion. LADICA will display the latest transcription of their conversation as it progresses. When the team concludes a section of their discussion, they can press “Extract key information” to identify the key points from their dialogue. If a key point relates to existing idea notes on screen, LADICA will also show the related note in each information card. Teams can then press on a key information card to create an idea note containing the same content.
@@ 4.3.3. Discussion reference layer
Establishing and preserving mutual awareness is essential for team coordination and recognizing opportunities for collaboration. To support this meta-cognitive process, LADICA is able to provide relation hints among ideas and update them periodically throughout the collaboration (DG4). As shown in Figure 9, users can press “Action” then “Toggle relation hints” in the workspace menu to activate the relation hint generation. In addition, users can toggle the “cross-user only” to only generate ideas created by different team members. To reflect the relations among the latest ideas, LADICA will update the relation hints every 10 seconds when the function is activated. Note that this feature is designed to indicate relationships to catalyze initial discussion rather than to suggest specific discussion content. To achieve this, we instruct the LLM to generate relation types based on the ConceptNet relation taxonomy (Speer et al., 2017), which encompasses a wide range of types and maintains a level of granularity that provides teams with ample room to explore the relationships.
@@
Our findings indicate that team members often struggle to keep the whiteboard content aligned with the dynamic flow of their discussions. To address this challenge, LADICA is designed to intelligently retrieve relevant existing idea nodes based on the content of the conversation (DG2). Users can activate this feature by pressing the “Start Recording” button at the beginning of their collaborative session (Figure 6). This prompts LADICA to start recording while simultaneously analyzing the ongoing discussion to identify mentions or references related to existing ideas. Upon ending the session with the ’Stop Recording’ button, LADICA processes the recorded conversation. The system then retrieves the existing ideas that have relevance score with its corresponding dialogue reference in the discussion over a predefined threshold (see Section 4.4.5). For each relevant ideas, LADICA also shows the corresponding dialogue reference.
@@
A crucial opportunity for generating new insights is to reflect on team’s ideas within group, which sparks diverse perspective in discussion. To facilitate this process, LADICA provides group-based discussion hint (Figure 10) that is able to generate hints of discussion topics about the theme of the group and existing ideas within the group. Teams can access this feature by pressing [Uncaptioned image] in the group sidebar (Figure 1, D). Users can provide further instructions for hint generation. It is important to note that the hints are not intended to generate ideas directly; instead, they are designed to inspire potential new directions for team thinking and discussion (DG4).